# Omegance: A Single Parameter for Various Granularities in Diffusion-Based Synthesis

### v1

@@ Abstract
In this work, we introduce a single parameter $\omega$, to effectively control granularity in diffusion-based synthesis. This parameter is incorporated during the denoising steps of the diffusion model’s reverse process. Our approach does not require model retraining, architectural modifications, or additional computational overhead during inference, yet enables precise control over the level of details in the generated outputs. Moreover, spatial masks or denoising schedules with varying $\omega$ values can be applied to achieve region-specific or timestep-specific granularity control. Prior knowledge of image composition from control signals or reference images further facilitates the creation of precise $\omega$ masks for granularity control on specific objects. To highlight the parameter’s role in controlling subtle detail variations, the technique is named Omegance, combining “omega” and “nuance”. Our method demonstrates impressive performance across various image and video synthesis tasks and is adaptable to advanced diffusion models. The code is available at https://github.com/itsmag11/Omegance.
@@ 1 Introduction
Diffusion models have emerged as powerful tools in image and art generation by progressively transforming random noise into coherent visual content through a learned iterative process. Diffusion models are particularly effective in generating high-quality, diverse results.
@@
Artists and designers often need to strategically decide where and how to apply details in their work. The level of detail in a piece of artwork or photograph can shape its visual harmony, order, and clarity, influencing how the viewer experiences and interprets it while guiding their focus [2]. The vanilla diffusion model does not inherently offer direct, fine-tuned control over the level of granularity in specific areas of an image. While the model can generate varying levels of detail across different images, its uniform generative process does not allow for easy manipulation of how much detail is rendered in different parts of the same image. The level of detail in an image can be challenging—or even impossible—to convey through text alone. For instance, reducing detail in the background while retaining high detail in the main subject (see the right case of Fig. 1(b) for illustration) is not straightforward.
@@
In this paper, we explore a novel yet ‘embarrassingly’ simple approach for controlling the level of detail in diffusion model outputs by scaling the predicted noise during each denoising step. Our method requires no modifications to the network architecture or timestep scheduling. Instead, we demonstrate that by dynamically adjusting the variance of the noise removed at each step, we can influence the granularity of the visual output. This simple yet flexible technique allows tailored adjustments in concept density and object texture, offering users a more nuanced control over the synthesized content.
@@
Our approach is appealing as the noise scaling is achieved with just a single parameter omega, $\omega$. Decreasing $\omega$ results in less noise being removed, leading the network to infer more complex scenes and richer textures. Conversely, increasing $\omega$ removes more noise, leading to smoother and simpler outputs. While applying our omega control globally over space and consistently over time can yield uniformly richer or smoother results, as shown in Fig. 1(a), more precise controls can be implemented both spatially and temporally. (1) Since the granularity requirement may vary within a single image, e.g., finer-grained details for areas requiring rich textures and complex visual elements, and coarser-grained details for areas demanding smooth transitions and high-level quality, we can use omega masks to customize the desired effects across different spatial regions. Examples of different spatial effects are shown in Fig. 1(b). The mask can be created either from user-provided strokes or generated using specific guiding conditions. (2) To better align with the diffusion denoising dynamics [10, 24], where the object shapes and image layouts typically emerge in the early stages and fine details in the later stages, we can implement omega schedules, adjusting the omega value over time for varying effects on layout and detailed textures. Examples are shown in Fig. 1(c).
@@
Our Omegance technique is not limited to any specific network architecture or denoising scheduler as long as the progressive diffusion denoising process is followed. Extensive experiments demonstrate Omegance’s ability to adapt to various diffusion-based synthesis tasks. Models evaluated include Stable Diffusion [16, 5] and FLUX [13] for text-to-image generation, SDEdit [15] and ControlNet [28] for image-to-image generation, SDXL-Inpainting [16] for image inpainting, ReNoise [7] for real-image editing, and Latte [14] and AnimateDiff [8] for text-to-video generation. Some examples are shown in Fig. 1. In all of the above applications, effective and smooth, nuanced control over the generated results is observed, demonstrating the effectiveness of our single-parameter granularity adjustment. In summary, our contributions are:
  • We propose a simple yet effective omega scaling technique that enables a single-parameter control over the granularity of generated content in diffusion-based models. Our method is training-free, incurs no additional inference cost, and is agnostic to network architectures.
  • We introduce both spatially adaptive omega masks and temporally dynamic omega schedules, which allow for localized and progressive control of granularity within generated images and videos.
  • We demonstrate the applicability of omega scaling on a vast range of image and video synthesis tasks.
@@ 2 Related Work
Diffusion-based Editing. Most previous diffusion-based editing methods focus on exploiting the visual-language association ability of CLIP [17] to edit visual content according to language guidance. Prompt-to-Prompt [9] and InstructPix2Pix [4] edit concepts in the output by modifying the cross-attention maps, which play a crucial role in aligning textual prompts with visual features during the generation process. SEGA [3] generates results following the semantic guidance of a target prompt during denoising. Wu et al. [25] find by mixing the text embedding of prompts with and without the target attribute, the output can preserve the original content while aligning with the desired attribute. Besides, SDEdit [15] adds noise to the modified image and utilizes diffusion prior to rationalize the edited parts as natural images. However, these previous methods are less effective when the editing target cannot be explicitly represented in language or denoted in the original image, and fail to provide a flexible way to edit the granularity of the output.
@@
Generation Quality Enhancement. Efforts have also been made to enhance the quality of content generated by diffusion models. Several works have explored modifications of Classifier-Free Guidance (CFG) to improve generation quality [11, 1, 20]. SAG [11] and PAG [1] substitute the null-text prediction in CFG with self-attention or perturbed self-attention maps, enabling high-quality, training- and condition-free generation. Sadat et al. [20] present a guidance strategy similar to CFG, applied between clean and perturbed text embeddings, to boost generation quality. While these methods effectively enhance quality globally, they lack the capability for spatially fine-grained control over details in the generated output. Another line of research leverages reinforcement learning from human feedback (RLHF) to fine-tune diffusion models for higher-quality results aligned with human preferences [26, 6, 27]. Xu et al. [26] present a general-purpose text-to-image human preference reward model and use it to fine-tune the diffusion model regarding human preference score. A similar approach is adopted in concurrent work DPOK [6]. Furthermore, Yang et al. [27] employ direct preference optimization (DPO), fine-tuning the diffusion model to align with human feedback without a separate reward model. While these methods produce outputs that reflect human preferences, they involve costly model fine-tuning and lack flexible control over output granularity. FreeU [22] was recently introduced to enhance the quality of diffusion model outputs, specifically targeting the U-Net architecture in the denoising process. This method involves jointly adjusting two scaling factors during inference: one for amplifying the backbone features and one for modulating the influence of the skip connections to better preserve details without over-smoothing or degrading high-frequency elements. While achieving noticeable quality improvement, FreeU is closely tied to the U-Net architecture and requires careful adjustment of its dual scaling parameters. In contrast, Omegance provides a simpler, more flexible, and architecture-agnostic approach to control the level of detail in diffusion models.
@@ 3.1 Diffusion Model Preliminaries
Diffusion models are powerful generative models that synthesize realistic images by iteratively predicting the noise added to a sample. They consist of two processes: In the forward process, Gaussian noise is progressively added to an initial latent $z_{0}$ that is directly decoded from $x_{0}$. Following Song et al. [23], we formulate the process as:
@@
where $z_{t}$ is the noisy latent at timestep $t$. The noise schedule $\alpha_{t}$ is defined as the cumulative product of $(1-\beta_{t})$: $\alpha_{t}=\prod_{i=1}^{t}(1-\beta_{i})$ where $\beta_{t}$ is a predefined variance schedule that controls the amount of noise added at each timestep [10]. In the reverse process, a pure Gaussian noise is transformed into coherent visual content through a learned denoising process. A general representation of the denoising process is:
@@
where $\delta_{t}$ and $\zeta_{t}$ are scaling factors of the current noisy signal and the noise prediction, respectively, that vary according to specific scheduler (see supp. for details), and $\epsilon_{\theta}(z_{t},t)$ is the noise prediction at time $t$ by the denoising network with parameters $\theta$. This formula characterizes the iterative denoising process, where each step aims to progressively move from a noisier latent $z_{t}$ towards clean $z_{0}$ to get a less noisy latent $z_{t-1}$.
@@
When $t=0$, $\mathrm{SNR}=1$ indicating pure image signal $z_{0}$. As $t$ increases, SNR decreases to 00, demonstrating pure noise $z_{T}$. During denoising, the model progressively aligns the SNR of each timestep with the SNR defined by the forward process. Since $\alpha_{t}$ is predefined by the noise schedule, the SNR in vanilla diffusion models remains fixed throughout the denoising process, limiting flexibility in controlling the amount of noise at each timestep.
@@
Figure 3: Global effects of Omegance. (Zoom-in for best view)
@@ 3.2 Omegance
Since the noise prediction in each denoising step is a standard Gaussian noise: $\epsilon_{\theta}(z_{t},t)\sim\mathcal{N}(0,1)$, multiplying it by a scaling factor $\omega$ preserves the mean at 0 while adjusting the variance proportionally to $\omega^{2}$. In practice, $\omega$ is rescaled to allow (,)($\infty$,$\infty$)( - ∞ , ∞ ) input range for finer-grained control and re-centered at 00 by:
$\omega=\mathcal{R}(\varpi)=L+\frac{U-L}{1+e^{-k\cdot\varpi}}$ (5)
Though it is a simple introduction of a coefficient to the denoising term, its influence on SNR and detail generation is worth investigating. Taking DDIM scheduler [23] as an example, the modified SNR during denoising is formulated as follows (see supp. for step-by-step derivations):
@@
However, in the flow-matching-based scheduler [5], the forward process learns a continuous transformation without the need for a stepwise noise addition schedule: $z_{t}=(1-t)z_{0}+t\epsilon$, where $\epsilon\sim\mathcal{N}(0,1)$, which slightly differs from Equ. (1). During the reverse process, the model predicts $v_{\theta}(z_{t},t)=\frac{dz_{t}}{dt}=\epsilon-z_{0}$, and moves one step forward with $z_{t-1}=z_{t}+dt\cdot v_{\theta}(z_{t},t)$. Here, the term $dt\cdot v_{\theta}(z_{t},t)$ represents the denoise amount as in the general formula Equ. (2), but is not necessarily a standard noise. To prevent mean-shifting, we apply an additional mean-preserving operation with Omegance in the flow matching scheduler.
@@
$\begin{split}m&=\mathbb{E}[dt\cdot v_{\theta}(z_{t},t)]\\ z_{t-1}^{\prime}&=z_{t}+[(dt\cdot v_{\theta}(z_{t},t)-m)\cdot\boldsymbol{\omega}+m]\end{split}$ (9)
Figure 4: Effects of Omegance in different denoising stages. The $\mathcal{S}^{1}(t)$ and $\mathcal{S}^{2}(t)$ schedules correspond to Early-Stage Enhancement (left) and Late-Stage Enhancement (right) cases in Fig. 1(c). The discrete step schedules are applied to ensure clearer effects on the image layout and fine-grained detail.
@@ 3.2.1 Omega Mask
The omega mask $\omega_{i,j}=\mathcal{M}(i,j)$ introduces a spatially varying control over the granularity within a single image by allowing different regions to have distinct $\omega$ values during the denoising process.
$\mathcal{M}$ is a mask $\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$ where $H^{\prime}=H/f,W^{\prime}=W/f$ are the original image dimension $H,W$ scaled by the VAE downsampling factor $f$. The mask can be obtained from user-provided strokes, segmentation masks, or automatically generated from control signals like pose skeleton, depth map, etc. in both discrete and continuous manners as illustrated in Fig. 8.
This spatial control leverages the locality of the denoising process, ensuring that adjustments to $\omega$ in one region do not affect the $\mathrm{SNR}$ or visual properties of neighboring areas. Such flexibility is valuable for applications requiring region-specific detail control within a single image, enabling fine-grained textures in focal regions while maintaining smoothness elsewhere.

### v2

@@ Abstract
In this work, we show that we only need a single parameter $\omega$ to effectively control granularity in diffusion-based synthesis. This parameter is incorporated during the denoising steps of the diffusion model’s reverse process. This simple approach does not require model retraining or architectural modifications and incurs negligible computational overhead, yet enables precise control over the level of details in the generated outputs. Moreover, spatial masks or denoising schedules with varying $\omega$ values can be applied to achieve region-specific or timestep-specific granularity control. External control signals or reference images can guide the creation of precise $\omega$ masks, allowing targeted granularity adjustments. Despite its simplicity, the method demonstrates impressive performance across various image and video synthesis tasks and is adaptable to advanced diffusion models. The code is available at https://github.com/itsmag11/Omegance.
@@ 1 Introduction
Diffusion models have emerged as powerful tools in image and art generation by progressively transforming random noise into coherent visual content through a learned iterative process [15, 30, 34, 27, 3, 16, 23]. They have become the dominant paradigm for high-quality image synthesis, offering strong diversity and controllability.
@@
Artists and designers often need to strategically decide where and how to apply details in their work. The level of detail in a piece of artwork or photograph can shape its visual harmony, order, and clarity, influencing how the viewer experiences and interprets it while guiding their focus [2, 7]. The vanilla diffusion model does not inherently offer direct, fine-tuned control over the level of granularity in specific areas of an image. While the model can generate varying levels of detail across different images, its uniform generative process does not allow for easy manipulation of how much detail is rendered in different parts of the same image. The level of detail in an image can be challenging—or even impossible—to convey through text alone. For instance, reducing detail in the background while retaining high detail in the main subject (see the right case of Fig. 1(b) for illustration) is not straightforward.
@@
In this paper, we explore a novel yet simple approach for controlling the level of detail in diffusion model outputs by scaling the predicted noise during each denoising step. The method does not require any network architecture or noise scheduling modifications. Instead, we demonstrate that it can influence the granularity of the visual output by dynamically adjusting the variance of the removed noise at each step. While variance scaling is a fundamental operation in diffusion models, to the best of our knowledge, it has not been systematically explored as a means for fine-grained granularity control. This simple yet flexible technique allows tailored adjustments in concept density and object texture, offering users a more nuanced control over the synthesized content.
@@
The presented approach is named Omegance, combining “omega” and “nuance”. It is appealing as it enables noise scaling with a single parameter, $\omega$. Decreasing $\omega$ results in less noise being removed, leading the network to infer more complex scenes and richer textures. Conversely, increasing $\omega$ removes more noise, leading to smoother and simpler outputs. While applying our omega control globally over space and consistently over time can yield uniformly richer or smoother results, as shown in Fig. 1(a), more precise controls can be implemented both spatially and temporally. (1) Since the granularity requirement may vary within a single image, e.g., finer-grained details for areas requiring rich textures and complex visual elements, and coarser-grained details for areas demanding smooth transitions and high-level quality, we can use omega masks to customize the desired effects across different spatial regions. Examples of different spatial effects are shown in Fig. 1(b). The mask can be created either from user-provided strokes or generated using specific guiding conditions. (2) To better align with the diffusion denoising dynamics [15, 40], where the object shapes and image layouts typically emerge in the early stages and fine details in the later stages, we can implement omega schedules, adjusting the omega value over time for varying effects on layout and detailed textures. Examples are shown in Fig. 1(c).
@@
Omegance is not limited to any specific network architecture or denoising scheduler as long as the progressive diffusion denoising process is followed. Extensive experiments demonstrate Omegance’s ability to adapt to various diffusion-based synthesis tasks. Models evaluated include Stable Diffusion [30, 8] and FLUX [22] for text-to-image generation, SDEdit [26] and ControlNet [48] for image-to-image generation, SDXL-Inpainting [30] for image inpainting, ReNoise [10] for real-image editing, and Latte [25] and AnimateDiff [11] for text-to-video generation. Some examples are shown in Fig. 1. In all of the above applications, effective and smooth, nuanced control over the generated results is observed, demonstrating the effectiveness of our single-parameter granularity adjustment.
@@ 2 Related Work
Diffusion-based Editing. Most previous diffusion-based editing methods focus on exploiting the visual-language association ability of CLIP [31] to edit visual content according to language guidance. Prompt-to-Prompt [12] and InstructPix2Pix [5] edit concepts in the output by modifying the cross-attention maps, which play a crucial role in aligning textual prompts with visual features during the generation process. SEGA [4] generates results following the semantic guidance of a target prompt during denoising. Wu et al. [45] find by mixing the text embedding of prompts with and without the target attribute, the output can preserve the original content while aligning with the desired attribute. Besides, SDEdit [26] adds noise to the modified image and utilizes diffusion prior to rationalize the edited parts as natural images. These methods struggle when the desired edits cannot be clearly described using language or inferred from the original image , and fail to provide a flexible way to edit the granularity of the output.
Generation Quality Enhancement. Efforts have also been made to enhance the quality of content generated by diffusion models. Several works have explored modifications of Classifier-Free Guidance (CFG) [14] to improve generation quality [17, 1, 33]. SAG [17] and PAG [1] substitute the null-text prediction in CFG with self-attention or perturbed self-attention maps, enabling high-quality, training- and condition-free generation. Sadat et al. [33] present a guidance strategy similar to CFG, applied between clean and perturbed text embeddings, to boost generation quality. While these methods effectively enhance quality globally, they lack the capability for spatially fine-grained control over details in the generated output [15, 39, 19, 28].
@@
Another line of research leverages reinforcement learning from human feedback (RLHF) [29, 32, 6, 36] to fine-tune diffusion models for higher-quality results aligned with human preferences [46, 9, 47]. Xu et al. [46] present a general-purpose text-to-image human preference reward model and use it to fine-tune the diffusion model regarding human preference score. A similar approach is adopted in concurrent work DPOK [9]. Furthermore, Yang et al. [47] employ direct preference optimization (DPO), fine-tuning the diffusion model to align with human feedback without a separate reward model. While these methods produce outputs that reflect human preferences, they involve costly model fine-tuning and lack flexible control over output granularity. FreeU [38] was recently introduced to enhance the quality of diffusion model outputs, specifically targeting the U-Net architecture in the denoising process. This method involves jointly adjusting two scaling factors during inference: one for amplifying the backbone features and one for modulating the influence of the skip connections to better preserve details without over-smoothing or degrading high-frequency elements. While achieving noticeable quality improvement, FreeU is closely tied to the U-Net architecture and requires careful adjustment of its dual scaling parameters. In contrast, Omegance provides a simpler, more flexible, and architecture-agnostic approach to control the level of detail in diffusion models.
Noise Scheduling. Noise scheduling is also crucial for diffusion model performance, impacting generation quality and stability. The linear and cosine schedulers [28] are widely used, with the linear scheduler applying uniform noise variance and the cosine scheduler preserving high-frequency details longer. Lin et al. [24] identify flaws for not enforcing zero terminal SNR in previous schedulers and propose a rescaled scheduler, correcting noise variance scaling for improved stability and fidelity. These findings underscore the need for careful noise schedule design to enhance sample quality. Unlike traditional noise schedulers that apply fixed, global denoising adjustments, which can be unpredictable and require model retraining, Omegance offers a lightweight, interpretable, and architecture-agnostic mechanism for both global and localized detail control, seamlessly integrating with existing schedulers.
@@ 3.1 Diffusion Model Preliminaries
Diffusion models are powerful generative models that synthesize realistic images by iteratively predicting the noise added to a sample. They consist of two processes: In the forward process, Gaussian noise $\epsilon$ is progressively added to an initial latent $z_{0}$ that is directly decoded from $x_{0}$. Following Song et al. [39], we formulate the process as:
@@
where $z_{t}$ is the noisy latent at timestep $t$. The noise schedule $\alpha_{t}$ is defined as the cumulative product of $(1-\beta_{t})$: $\alpha_{t}=\prod_{i=1}^{t}(1-\beta_{i})$ where $\beta_{t}$ is a predefined variance schedule that controls the amount of noise added at each timestep [15].  [^1] Note that some papers [15, 28, 24] denote $(1-\beta_{t})$ as $\alpha_{t}$ and $\prod_{i=1}^{t}\alpha_{t}$ as $\bar{\alpha_{t}}$. In the reverse process, pure Gaussian noise is transformed into coherent visual content through a learned denoising process. A general representation of the denoising process is:
@@
where $\delta_{t}$ and $\zeta_{t}$ are scaling factors of the current noisy signal and the noise prediction, respectively, that vary according to specific scheduler (see supp. Sec.A for details), and $\epsilon_{\theta}(z_{t},t)$ is the noise prediction at time $t$ by the denoising network with parameters $\theta$. This formula characterizes the iterative denoising process, where each step aims to progressively move from a noisier latent $z_{t}$ towards clean $z_{0}$ to get a less noisy latent $z_{t-1}$.
@@
When $t\rightarrow 0$, $\mathrm{SNR}\rightarrow\infty$ indicating pure image signal $z_{0}$. As $t$ increases, SNR decreases to 00, demonstrating pure noise $z_{T}$. During denoising, the model progressively aligns the SNR of each timestep with the SNR defined by the forward process. Since $\alpha_{t}$ is predefined by the noise schedule, the SNR in vanilla diffusion models remains fixed throughout the denoising process, limiting flexibility in controlling the amount of noise at each timestep.
@@
Figure 3: Global effects of Omegance. The models indicated below are the base models. The middle row shows the original base model results. The top and bottom rows are Omegance results with detail enhancement and suppression, respectively. Omegance can effectively add or remove details without harming visual quality or modify the entire image completely, making it a flexible tool for practical use.
@@ 3.2 Omegance
Since the diffusion model is trained to predict Gaussian noise, $\epsilon_{\theta}(z_{t},t)$ can be viewed as an estimation of the standard Gaussian noise $\epsilon\sim\mathcal{N}(0,1)$. While $\epsilon_{\theta}$ itself is a deterministic output and not necessarily Gaussian-distributed, scaling it by a factor $\omega$ still serves to control the effective denoising direction during sampling. This controls how much detail is recovered, without altering the underlying direction toward the clean signal. In practice, $\omega$ is rescaled by $\omega=\mathcal{R}(\varpi)$ to allow input $\varpi\in(-\infty,\infty)$ for finer-grained control and re-centered at 00 (see formulation and sensitivity test in supp. Sec.J.3).
Though it is a simple introduction of a coefficient to the denoising term, its influence on SNR and detail generation is worth investigating. Taking DDIM scheduler [39] as an example, the modified SNR during denoising is formulated as follows (see supp. Sec.C for step-by-step derivations):
@@
However, in the flow-matching-based scheduler [8], the forward process learns a continuous transformation without the need for a stepwise noise addition schedule: $z_{t}=(1-t)z_{0}+t\epsilon$, where $\epsilon\sim\mathcal{N}(0,1)$, which slightly differs from Equ. (1). During the reverse process, the model predicts $v_{\theta}(z_{t},t)=\frac{dz_{t}}{dt}=\epsilon-z_{0}$, and moves one step forward with $z_{t-1}=z_{t}+dt\cdot v_{\theta}(z_{t},t)$. Here, the term $dt\cdot v_{\theta}(z_{t},t)$ represents the denoise amount as in the general formula Equ. (2), but is not necessarily a standard noise. To prevent mean-shifting which causes unwanted color change (details in supp. Sec.I.2), we apply a mean-preserving operation with Omegance in the flow matching scheduler.
@@
$\begin{split}m&=\mathbb{E}[dt\cdot v_{\theta}(z_{t},t)]\\ z_{t-dt}^{\prime}&=z_{t}+[(dt\cdot v_{\theta}(z_{t},t)-m)\cdot\boldsymbol{\omega}+m]\end{split}$ (8)
Figure 4: Illustration of $\omega$ effect during denosing process. During the early stage ($t\in[T,\tau]$), a higher $\omega$ reduces layout complexity (blue region), while a lower $\omega$ enhances it (red region). In the late stage ($t\in[\tau,0]$), a higher $\omega$ suppresses fine-grained details, whereas a lower $\omega$ enhances them. The $\mathcal{S}^{1}(t)$ and $\mathcal{S}^{2}(t)$ schedules correspond to Early-Stage Enhancement (left) and Late-Stage Enhancement (right) cases in Fig. 1(c). More examples in Fig. 6.
@@ 3.2.1 Omega Mask
The omega mask $\omega_{i,j}=\mathcal{M}(i,j)$ introduces a spatially varying control over the granularity within a single image by allowing different regions to have distinct $\omega$ values during the denoising process. $\mathcal{M}$ is a mask $\in\mathbb{R}^{H^{\prime}\times W^{\prime}}$ where $H^{\prime}=H/f,W^{\prime}=W/f$ are the original image dimension $H,W$ scaled by the VAE downsampling factor $f$. The mask can be obtained from user-provided strokes, segmentation masks, or automatically generated from control signals like pose skeleton, depth map, etc. in both discrete and continuous manners as illustrated in Fig. 7. This spatial control leverages the locality of the denoising process, ensuring that adjustments to $\omega$ in one region do not affect the $\mathrm{SNR}$ or visual properties of neighboring areas. Such flexibility is valuable for applications requiring region-specific detail control within a single image, enabling fine-grained textures in focal regions while maintaining smoothness elsewhere.